# LLM-Driven Usefulness Labeling for IR Evaluation

### v1

@@ 5.3. RQ3: Ablation Study
To address our RQ3, we conduct an ablation study of the metrics provided to the LLM. We categorize the features into five levels: query features (Q), document features (D), relevance metrics (R), satisfaction metrics (S), and user behavior features (U) as shown in Table 4. For the ablation study, we keep the query (Q) and document features (D) constant while varying the relevance features (R), satisfaction features (S), and user behavior features (U). This approach aims to identify the most critical features required for pre-trained LLMs to effectively generate usefulness labels. In real-world applications using all possible inputs in the LLM prompt is token-intensive and cost-prohibitive. That’s why it is important to optimize feature selection to address these challenges. For conducting the ablation study, we chose Llama 3.3 70B-Instruct and GPT-4omini on the THUIR-KDD’19 session data as these two models were consistently outperforming other models in usefulness judgment. From the ablation study shown in Table 4, observing the spearman’s correlation (), it is evident that the models seem to perform the best when all three features such as relevance (R), satisfaction (S) and user features (U) are provided along with query (Q) and document features (D) with a moderate positive correlation ( = 0.36 and 0.37). The next is followed by relevance (R) and satisfaction features (S) that also has a moderate positive correlation ( = 0.35). When satisfaction (S) and user features (U) are used individually it has the lowest correlation ( = 0.27 and 0.28). The findings of the ablation study align closely with our research questions and provide valuable insights into the role of specific features. Labels such as relevance and satisfaction significantly enhance an LLM’s ability to capture the nuanced contextual information within search session data. However, when features are provided independently, the LLM struggles to effectively evaluate usefulness, making it challenging to perform an accurate usefulness evaluation. This highlights the importance of combining relevant features to enable the LLM to interpret and evaluate search sessions effectively.
@@ Acknowledgements.
This work was supported in part by the cloud computing credits made available by the University of Washington’s InfoSeeking Lab in partnership with Amazon Web Services (AWS).References
@@
- Voorhees and Harman (2000) Ellen M Voorhees and Donna Harman. 2000. Overview of the sixth text retrieval conference (TREC-6). Information Processing & Management 36, 1 (2000), 3–35.

### v2

@@ 5.3. RQ3: Ablation Study
To address our RQ3, we conduct an ablation study of metrics provided to the LLM. We categorize the features into five levels: query features (Q), document features (D), relevance metrics (R), satisfaction metrics (S), and user behavior features (U) as shown in Table 4. For the ablation study, we keep Q and D constant while changing R, S and U. This approach aims to identify the most critical features required for LLMs to effectively generate usefulness labels. In real-world applications using all possible inputs in the LLM prompt is token-intensive and cost-prohibitive that’s why it is important to optimize feature selection. For conducting the ablation study, we chose Llama 3.3 70B-Instruct and GPT-4omini on the THUIR-KDD’19 session data as these two consistently outperformed others in usefulness judgment. From the ablation study shown in Table 4, observing the Spearman’s correlation () models seem to perform the best when all three features such as relevance (R), satisfaction (S) and user features (U) are provided along with query (Q) and document features (D) with a moderate positive correlation ( = 0.36 and 0.37). The next is followed by relevance (R) and satisfaction features (S) that also has a moderate positive correlation ( = 0.35). This shows that LLMs can infer usefulness if both relevance and satisfaction levels are present supporting prior research on usefulness (Mao et al., 2016b). Notably, when satisfaction (S) and user features (U) are used individually it shows the lowest correlation ( = 0.27 and 0.28) indicating that both features are associated with usefulness should be considered for improved performance. The findings of the ablation study highlights relevance and satisfaction as key in determining usefulness and both combined can improve LLM’s ability to capture nuanced contextual information. However, when features are provided independently, LLM struggles to effectively evaluate usefulness making it challenging to perform an accurate usefulness evaluation. This highlights the importance of combining features and focusing on relevance to enable the LLM to interpret and evaluate the usefulness of search sessions effectively.
A comparison between human annotators and LLM-based usefulness labels.
@@ Acknowledgements.
This work was partially supported by cloud computing credits provided through an Amazon Research Award to Dr. Chirag Shah.References
@@
- Fox et al. (2005) Steve Fox, Kuldeep Karnawat, Mark Mydland, Susan Dumais, and Thomas White. 2005. Evaluating implicit measures to improve web search. ACM Transactions on Information Systems (TOIS) 23, 2 (2005), 147–168.